# Excitation Spectrum and Spin Hamiltonian of the Frustrated Quantum Ising Magnet PrBWO

### v1

@@ Abstract
We present a thorough experimental investigation on single crystals of the rare-earth based frustrated quantum antiferromagnet PrBWO, a purported spin-liquid candidate on the breathing kagome lattice. Despite its disordered ground state and unusual excitation spectrum involving a coexistence of sharp spin-waves and broad continuum excitations, we show through a combination of thermodynamic, magnetometric, neutron- and muon spectroscopy experiments with detailed theoretical modeling that this material should be understood in a completely different framework. The crystal field splits the lowest quasi-doublet states into two singlets moderately coupled through frustrated superexchange, resulting in a simple effective Hamiltonian of an Ising model in a transverse magnetic field. While our neutron spectroscopy data do point to significant correlations within the kagome planes, the dominant interactions are out-of-plane, forming frustrated triangular spin-tubes through two competing ferro-antiferromagnetic bonds. The resulting ground state is a simple quantum paramagnet, but with significant modifications to both thermodynamic and dynamic properties due to small perturbations to the transverse field Ising model in the form of hyperfine enhanced nuclear moments and weak structural disorder.
@@ I Introduction
Case in point, non-Kramers ions in a low symmetry environment typically posses a quasi-doublet ground state, the physics of which can be mapped to an Ising model in a transverse magnetic field [wangCollectiveExcitationsMagnetic1968, chenIntrinsicTransverseField2019]. The transverse field Ising model (TFIM) is an archetype of quantum magnetism with only a handful of experimental realizations [sachdevQuantumPhaseTransitions2011a, coldeaQuantumCriticalityIsing2010]. Especially on geometrically frustrated lattices this model still holds tremendous promise, where the interplay of quantum fluctuations and extensive degeneracy promotes a plethora of new exotic phases [chenIntrinsicTransverseField2019, moessnerTwoDimensionalPeriodicFrustrated2000, moessnerIsingModelsQuantum2001a]. Intrinsic realizations of the TFIM Hamiltonian have gained increasing traction in recent years, with hallmark examples such as the triangular lattice material TmMgGaO [liPartialUpUpDownOrder2020, dunNeutronScatteringInvestigation2021] or the tripod kagome magnet HoMgSbO [dunQuantumClassicalSpin2020]. Nevertheless, good prototype materials are still relatively scarce and their experimental exploration has only just begun.
@@
The recently discovered family BWO of quantum antiferromagnets poses a promising route to realizing such physics [ashtarNewFamilyDisorderFree2020]. Trivalent rare-earth ions are arranged on a breathing kagome lattice in the hexagonal basal plane and the common problem of antisite disorder is suppressed by a large difference between ionic radii. All members investigated so far exhibit a significant frustration ratio [flavianMagneticPhaseDiagram2023a, zengLocalEvidenceCollective2021], although our recent work points to the potential relevance of significant inter-plane correlations between the kagome layers [flavianMagneticPhaseDiagram2023a].
@@
The compound PrBWO seems particularly interesting in this context. It retains the largest Weiss temperature K in the family [ashtarNewFamilyDisorderFree2020] and shows no signs of long-range order down to 0.1 K [zengLocalEvidenceCollective2021], while the low-symmetry environment of the non-Kramers Pr ion should allow for TFIM physics. Still, aside from a recent NMR study [zengLocalEvidenceCollective2021] reporting persistent collective fluctuations with a small field-dependent excitation gap, not much is known about the physics of this material.
@@
In this study, we present a thorough investigation of the magnetic ground state properties and excitation spectrum on single crystals of PrBWO. We utilize bulk magnetometry and inelastic neutron scattering on powders to characterize the single-ion ground state, showing that this compound can indeed be understood in the context of a frustrated TFIM Hamiltonian. Heat capacity, ultrasound and muon spin relaxation measurements are employed to map out the phase diagram, confirming a quantum disordered ground state in the entire phase space. Furthermore, the excitation spectrum is investigated through time of flight neutron spectroscopy on single crystals, revealing a coexistence of sharp spin-waves and broad continuum excitations. We propose a minimal model Hamiltonian capable of semi-quantitatively capturing all features of the spectrum, explaining these observations in terms of dispersive crystal field excitons partially broadened by weak structural disorder. A dominant dispersion perpendicular to the purported kagome planes confirms that - just like NdBWO [flavianMagneticPhaseDiagram2023a] - this system is not a pure kagome magnet, but realizes a complex exchange topology of twisted triangular spin-tubes, moderately coupled by the frustrated planar interactions. The source of chemical disorder is investigated with powder diffraction and magnetometry measurements, pointing to a scenario where of Pr impurities and a concomitant excess of oxygen create a distribution of crystal field splittings, with drastic effects on the thermodynamic and magnetic properties. Taken together, PrBWO turns out to be a relatively simple quantum paramagnet with a singlet ground state. Nevertheless, the interplay between frustrated Ising-like magnetic moments, local distortions brought about by structural disorder and strongly hyperfine-coupled nuclear spins gives rise to a number of unusual modifications to this fascinating model system.
@@ II.1 The Material
| Bond Distance (Å) | Bond Angle () | ||||||
3.96 |
2.43 / 2.57 |
2.51 / 2.57 |
104.9 |
102.5 |
3.96 |
2.43 / 2.63 |
2.34 / 2.37 |
103.0 |
114.5 |
4.25 |
2.46 / 2.57 |
115.1 |
4.39 |
2.46 / 2.51 |
124.4 |
4.94 |
2.57 / 2.63 |
143.4 |
PrBWO crystalizes in a hexagonal structure with space group and lattice parameters and . Its magnetism stems from the six Pr ions in each unit cell with electronic configuration. Due to the trivial site symmetry (point group ), the non-Kramers multiplet is fully split by the crystal electric field (CEF) and the single-ion ground state must be a singlet. Any potential low-energy magnetic properties then result from a mixing with low-lying CEF excitations through perturbations such as superexchange. If these interactions are sufficiently strong and not exceedingly frustrated, the lowest excitation branch may condense into a soft mode, giving rise to long-range magnetic order [thalmeierInducedQuantumMagnetism2023].
As shown in Figure 1(a), the Pr ions are arranged into a breathing-kagome lattice in the crystallographic basal plane, stacked in an AB alternating fashion. Neighboring triangles have slightly different sizes, resulting in a breathing anisotropy with inequivalent in-plane exchanges and . However, the shortest Pr-O-Pr bond actually corresponds to an out of plane coupling connecting the larger triangles and has to be considered on equal footing with the kagome interactions [flavianMagneticPhaseDiagram2023a] (see Table 1 for details). Due to several inequivalent superexchange pathways, this nearest-neighbor interaction must be divided into two seperate couplings and . These can partially release the frustration, making up twisted triangular spin-tubes along the crystallographic -axis as schematically depicted in Figure 1(c), with () chains running (counter-)clockwise. While the kagome interactions must be antiferromagnetic following the Kanamori-Goodenough rules [goodenoughMagnetismChemicalBond], the smaller bond angles for nearest neighbors could result in either ferromagnetic (FM) or antiferromagnetic (AFM) correlations.
@@ II.2 Experimental Details
Powder samples of PrBWO were prepared by solid-state reaction. High quality single-crystals were grown from these powders using a flux method analogous to [majchrowskiGrowthSpectroscopicCharacterization2003], resulting in needle-shaped green transparent crystals with six hexagonal faces and a typical mass of order mg, see Figure 2(a). B-enriched samples for neutron studies, as well as non-magnetic LaBWO powders were prepared by the same methods. The structure and sample quality of all powders/single crystals was confirmed through x-ray diffraction, using Rigaku MiniFlex and Bruker APEX-II instruments respectively. Furthermore, a low temperature structural refinement was performed at the thermal neutron diffractometer HRPT [FisherPhysBCM2000] at Paul Scherrer Institute (PSI, Switzerland). 10 g of PrBWO powders were sealed in a vanadium can and diffractograms were collected at K for two different wavelengths and .
@@
Magnetic susceptibility measurements on a 3.6 mg single crystal were carried out using an Magnetic Property Measurement System (MPMS) SQUID magnetometer in a probing field T along the crystallographic axes a, b and c. The low-temperature magnetization was measured up to 7 T in the same experimental configuration. High-field magnetization curves were collected with a compensated pickup coil setup in pulsed fields up to 50 T at the High Magnetic Field Laboratory in Dresden [SkourskiPRB832011]; the signal was calibrated to absolute units using the low-field SQUID data. Heat capacity of a 0.16 mg single crystal sample was measured by means of relaxation calorimetry on a 14 T Physical Property Measurement System (PPMS) with a He-He dilution refrigerator insert. Sound velocity measurements of the longitudinal mode (polarization and propagation along the crystallographic c-axis) were carried out at the Dresden High Magnetic Field Laboratory using a pulse-echo method with phase-sensitive detection [luthiPhysicalAcousticsSolid1967, WolfPhysBCM2001]. 36 Y-cut LiNbO transducers were bonded to two parallel faces of a mm long single cystal sample. Measurements were taken in the fundamental MHz mode using a He-He dilution system in static fields up to 18 T, as well as a He cryostat in pulsed fields up to 60 T.
@@
Muon spin relaxation (SR) measurements on powder samples were carried out on the newly commissioned FLAME spectrometer at PSI. A pressed powder pellet was mounted on a Cu sample holder and installed in a He-He dilution refrigerator. Positively charged muons were implanted in the sample and asymmetry profiles of the subsequent muon decay were registered in the forward- and backward positron detectors. Measurements were taken in the temperature range 30 mK 200 K and in fields up to 0.5 T. The SR spectra were analyzed using the MUSRFIT [suterMusrfitFreePlatformIndependent2012] software package.
@@
Inelastic neutron scattering (INS) experiments on powders were performed at the thermal neutron triple-axis-spectrometer EIGER [StuhrPhysBCM2017] at PSI. 12 g of PrBWO were sealed in an aluminum can and installed in a He orange cryostat. We chose a final wavelength of = 2.66 Å ( = 2.36 Å), using a pyrolytic graphite filter to minimize higher-harmonic scattering. Energy scans at constant scattering angle were taken at various temperatures. Neutron spectroscopy measurements on single crystals were carried out at the cold disc-chopper spectrometer AMATERAS [NakajimaJPSJS802011] in the Japan Proton Accelerator Research Complex (J-PARC). 282 hexagonal crystals were coaligned in the scattering plane for a total sample mass of 410 mg, depicted in Figure 2(b). The probe was installed in a He cryostat and measurements were carried out at base temperature K with incident energies meV and 2.63 meV, providing coverage of the entire first Brillouin zone at an elastic FWHM energy resolution of eV and 45 eV respectively. The resulting spectra were analyzed using the Horace [ewingsHORACESoftwareAnalysis2016] and Sunny [sunny_git] software packages.
@@ III.1 Magnetometry
-7.74 |
2.76 |
-7.98 |
2.80 |
-5.72 |
3.71 |
The bulk magnetic susceptibility along different crystallographic axes is shown in Figure 3(a). There are no signs of magnetic order down to 2 K and a simple fit to the inverse susceptibilities in the range K results in an average Weiss temperature K (see Table 2 for details), in agreement with previous reports [ashtarNewFamilyDisorderFree2020]. As the temperature is lowered, a pronounced easy axis anisotropy emerges along the -direction, while the planar response remains nearly axially symmetric over the entire temperature range. This should not be related directly to the single-ion anisotropy of Pr though, as the bulk susceptibility measured in experiment represents an average over the six ions in the unit cell, whose easy axes are related by rotations in the hexagonal plane (see Figure 4(c)). In the low temperature limit, the susceptibility exhibits a small downturn, which could result either from interactions or from a saturation of the large Van-Vleck susceptibility expected for a singlet ground state system.
@@ III.2 CEF Spectrum
The inelastic neutron spectra measured on B-substituted powders of PrBWO are shown in Figure 4(a). Apart from the strong quasielastic scattering, the only prominent feature is an intense peak at meV. This mode nicely matches the energy scale seen in magnetometry. It softens slightly with increasing temperature due to thermal expansion and its integrated intensity closely follows the Debye-Waller factor, confirming it is of magnetic origin. Although additional features are observed at higher energies meV (not shown), they are all shown to be either spurious scattering in the monochromator/analyzer or phonons.
@@ III.3 Single-Ion Ground State
where are the Stevens operators [stevensMatrixElementsOperator1952, hutchingsPointChargeCalculationsEnergy1964] and is the free-ion Landé factor. However, due to the trivial magnetic point symmetry, there are 27 allowed crystal field parameters , making an unrestricted fit of experimental data to this Hamiltonian severely underconstrained. Instead, we rely on an effective point-charge model [scheiePyCrystalFieldSoftwareCalculation2021, dunEffectivePointchargeAnalysis2021, hutchingsPointChargeCalculationsEnergy1964]: An estimate of the CEF parameters is obtained directly from the experimentally determined crystal structure (see Sec. III.9) by calculating the electrostatic repulsion of neighboring ligands. We can diagonalize the resulting Hamiltonian to obtain the full spectrum and single-ion wavefunctions, allowing a direct comparison with experiment. By varying the effective charges of the three inequivalent O positions as well as one mean field exchange parameter in a simultaneous fit to INS, Magnetization and Susceptibility data, we obtain an accurate approximation of the single-ion Hamiltonian that reproduces all experimental results.
@@
The local frame of each Pr ion is tilted by from the crystallographic -axis (see Figure 1(b)). In this reference frame, the quasi-doublet wavefunctions are well approximated by the (anti-)symmetric superpositions
The single-ion properties are axially symmetric, with perfect Ising anisotropy and due to the non-Kramers nature, while over 95% of the weights are concentrated on the maximal components. The weakened anisotropy observed in bulk arises naturally from averaging over the 6 equivalent Pr ions in the unit cell. A sketch of the ground state anisotropy is depicted in Figure 4(c), illustrating the probability to observe a spin with any given orientation.

### v2

# Excitation Spectrum and Spin Hamiltonian of the Frustrated Quantum Ising Magnet Pr3BWO9

@@ Abstract
We present a thorough experimental investigation on single crystals of the rare-earth based frustrated quantum antiferromagnet Pr3BWO9, a purported spin-liquid candidate on the breathing kagome lattice. This material possesses a disordered ground state with an unusual excitation spectrum involving a coexistence of sharp spin-waves and broad continuum excitations. Nevertheless, we show through a combination of thermodynamic, magnetometric and spectroscopic probes with detailed theoretical modeling that it should be understood in a completely different framework. The crystal field splits the lowest quasi-doublet states into two singlets moderately coupled through frustrated superexchange, resulting in a simple effective Hamiltonian of an Ising model in a transverse magnetic field. While our neutron spectroscopy data do point to significant correlations within the kagome planes, the dominant interactions are out-of-plane, forming frustrated triangular spin-tubes through two competing ferro-antiferromagnetic bonds. The resulting ground state is a simple quantum paramagnet, but with significant modifications to both thermodynamic and dynamic properties due to small perturbations to the transverse field Ising model in the form of hyperfine enhanced nuclear moments and weak structural disorder.
@@ I Introduction
Case in point, non-Kramers ions in a low symmetry environment typically posses a quasi-doublet ground state, the physics of which can be mapped to an Ising model in a transverse magnetic field [wangCollectiveExcitationsMagnetic1968, chenIntrinsicTransverseField2019]. The transverse field Ising model (TFIM) is an archetype of quantum magnetism with only a handful of experimental realizations [sachdevQuantumPhaseTransitions2011a, coldeaQuantumCriticalityIsing2010]. Especially on geometrically frustrated lattices this model still holds tremendous promise, where the interplay of quantum fluctuations and extensive degeneracy promotes a plethora of new exotic phases [chenIntrinsicTransverseField2019, moessnerTwoDimensionalPeriodicFrustrated2000, moessnerIsingModelsQuantum2001a]. Intrinsic realizations of the TFIM Hamiltonian have gained increasing traction in recent years, with hallmark examples such as the triangular lattice material TmMgGaO4 [liPartialUpUpDownOrder2020, dunNeutronScatteringInvestigation2021] or the tripod kagome magnet Ho3Mg2Sb3O14 [dunQuantumClassicalSpin2020]. Nevertheless, good prototype materials are still relatively scarce and their experimental exploration has only just begun.
@@
The recently discovered family BWO9 of quantum antiferromagnets poses a promising route to realizing such physics [ashtarNewFamilyDisorderFree2020]. Trivalent rare-earth ions are arranged on a breathing kagome lattice in the hexagonal basal plane and the common problem of antisite disorder is suppressed by a large difference between ionic radii. All members investigated so far exhibit a significant frustration ratio [flavianMagneticPhaseDiagram2023a, zengLocalEvidenceCollective2021], although our recent work points to the potential relevance of significant inter-plane correlations between the kagome layers [flavianMagneticPhaseDiagram2023a].
@@
The compound Pr3BWO9 seems particularly interesting in this context. It retains the largest Weiss temperature K in the family [ashtarNewFamilyDisorderFree2020] and shows no signs of long-range order down to 0.1 K [zengLocalEvidenceCollective2021], while the low-symmetry environment of the non-Kramers Pr3+ ion should allow for TFIM physics. Still, aside from a recent NMR study [zengLocalEvidenceCollective2021] reporting persistent collective fluctuations with a small field-dependent excitation gap, not much is known about the physics of this material.
@@
In this study, we present a thorough investigation of the magnetic ground state properties and excitation spectrum on single crystals of Pr3BWO9. We utilize bulk magnetometry and inelastic neutron scattering on powders to characterize the single-ion ground state, showing that this compound can indeed be understood in the context of a frustrated TFIM Hamiltonian. Heat capacity, ultrasound and muon spin relaxation measurements are employed to map out the phase diagram, confirming a quantum disordered ground state in the entire phase space. Furthermore, the excitation spectrum is investigated through time of flight neutron spectroscopy on single crystals, revealing a coexistence of sharp spin-waves and broad continuum excitations. We propose a minimal model Hamiltonian capable of semi-quantitatively capturing all features of the spectrum, explaining these observations in terms of dispersive crystal field excitons partially broadened by weak structural disorder. A dominant dispersion perpendicular to the purported kagome planes confirms that - just like Nd3BWO9 [flavianMagneticPhaseDiagram2023a] - this system is not a pure kagome magnet, but realizes a complex exchange topology of twisted triangular spin-tubes, moderately coupled by the frustrated planar interactions. The source of chemical disorder is investigated with powder diffraction and magnetometry measurements, pointing to a scenario where of Pr4+ impurities and a concomitant excess of oxygen create a distribution of crystal field splittings, with drastic effects on the thermodynamic and magnetic properties. Taken together, Pr3BWO9 turns out to be a relatively simple quantum paramagnet with a singlet ground state. Nevertheless, the interplay between frustrated Ising-like magnetic moments, local distortions brought about by structural disorder and strongly hyperfine-coupled nuclear spins gives rise to a number of unusual modifications to this fascinating model system.
@@ II.1 The Material
| Bond Distance (Å) | Bond Angle (∘) | ||||||||
| 3.96 | 2.43 / 2.57 | 2.51 / 2.57 | - | 104.9 | 102.5 | - | |||
| 3.96 | 2.43 / 2.63 | - | 2.34 / 2.37 | 103.0 | - | 114.5 | |||
| 4.25 | - | 2.46 / 2.57 | - | - | 115.1 | - | |||
| 4.39 | - | 2.46 / 2.51 | - | - | 124.4 | - | |||
| 4.94 | 2.57 / 2.63 | - | - | 143.4 | - | - | |||
Pr3BWO9 crystalizes in a hexagonal structure with space group and lattice parameters and . Its magnetism stems from the six Pr3+ ions in each unit cell with electronic configuration. Due to the trivial site symmetry (point group ), the non-Kramers multiplet is fully split by the crystal electric field (CEF) and the single-ion ground state must be a singlet. Any potential low-energy magnetic properties then result from a mixing with low-lying CEF excitations through perturbations such as superexchange. If these interactions are sufficiently strong and not exceedingly frustrated, the lowest excitation branch may condense into a soft mode, giving rise to long-range magnetic order [thalmeierInducedQuantumMagnetism2023].
As shown in Figure 1(a), the Pr3+ ions are arranged into a breathing-kagome lattice in the crystallographic basal plane, stacked in an AB alternating fashion. Neighboring triangles have slightly different sizes, resulting in a breathing anisotropy with inequivalent in-plane exchanges and . However, the shortest Pr-O-Pr bond actually corresponds to an out of plane coupling connecting the larger triangles and has to be considered on equal footing with the kagome interactions [flavianMagneticPhaseDiagram2023a] (see Table 1 for details). Due to several inequivalent superexchange pathways, this nearest-neighbor interaction must be divided into two seperate couplings and . These can partially release the frustration, making up twisted triangular spin-tubes along the crystallographic -axis as schematically depicted in Figure 1(c), with () chains running (counter-)clockwise. While the kagome interactions must be antiferromagnetic following the Kanamori-Goodenough rules [goodenoughMagnetismChemicalBond], the smaller bond angles for nearest neighbors could result in either ferromagnetic (FM) or antiferromagnetic (AFM) correlations.
@@ II.2 Experimental Details
Powder samples of Pr3BWO9 were prepared by solid-state reaction. High quality single-crystals were grown from these powders using a flux method analogous to [majchrowskiGrowthSpectroscopicCharacterization2003], resulting in needle-shaped green transparent crystals with six hexagonal faces and a typical mass of order mg, see Figure 2(a). 11B-enriched samples for neutron studies, as well as non-magnetic La3BWO9 powders were prepared by the same methods. The structure and sample quality of all powders/single crystals was confirmed through x-ray diffraction, using Rigaku MiniFlex and Bruker APEX-II instruments respectively. Furthermore, a low temperature structural refinement was performed at the thermal neutron diffractometer HRPT [FisherPhysBCM2000] at Paul Scherrer Institute (PSI, Switzerland). 10 g of PrBWO9 powders were sealed in a vanadium can and diffractograms were collected at K for two different wavelengths and .
@@
Magnetic susceptibility measurements on a 3.6 mg single crystal were carried out using an Magnetic Property Measurement System (MPMS) SQUID magnetometer in a probing field T along the crystallographic axes a∗, b and c. The low-temperature magnetization was measured up to 7 T in the same experimental configuration. High-field magnetization curves were collected with a compensated pickup coil setup in pulsed fields up to 50 T at the High Magnetic Field Laboratory in Dresden [SkourskiPRB832011]; the signal was calibrated to absolute units using the low-field SQUID data. Heat capacity of a 0.16 mg single crystal sample was measured by means of relaxation calorimetry on a 14 T Physical Property Measurement System (PPMS) with a 3He-4He dilution refrigerator insert. Sound velocity measurements of the longitudinal mode (polarization and propagation along the crystallographic c-axis) were carried out at the Dresden High Magnetic Field Laboratory using a pulse-echo method with phase-sensitive detection [luthiPhysicalAcousticsSolid1967, WolfPhysBCM2001]. 36∘ Y-cut LiNbO3 transducers were bonded to two parallel faces of a mm long single cystal sample. Measurements were taken in the fundamental MHz mode using a 3He-4He dilution system in static fields up to 18 T, as well as a 3He cryostat in pulsed fields up to 60 T.
@@
Muon spin relaxation (SR) measurements on powder samples were carried out on the newly commissioned FLAME spectrometer at PSI. A pressed powder pellet was mounted on a Cu sample holder and installed in a 3He-4He dilution refrigerator. Positively charged muons were implanted in the sample and asymmetry profiles of the subsequent muon decay were registered in the forward- and backward positron detectors. Measurements were taken in the temperature range 30 mK 200 K and in fields up to 0.5 T. The SR spectra were analyzed using the MUSRFIT [suterMusrfitFreePlatformIndependent2012] software package.
@@
Inelastic neutron scattering (INS) experiments on powders were performed at the thermal neutron triple-axis-spectrometer EIGER [StuhrPhysBCM2017] at PSI. 12 g of PrBWO9 were sealed in an aluminum can and installed in a 4He orange cryostat. We chose a final wavelength of = 2.66 Å-1 ( = 2.36 Å), using a pyrolytic graphite filter to minimize higher-harmonic scattering. Energy scans at constant scattering angle were taken at various temperatures. Neutron spectroscopy measurements on single crystals were carried out at the cold disc-chopper spectrometer AMATERAS [NakajimaJPSJS802011] in the Japan Proton Accelerator Research Complex (J-PARC). 282 hexagonal crystals were coaligned in the scattering plane for a total sample mass of 410 mg, depicted in Figure 2(b). The probe was installed in a 3He cryostat and measurements were carried out at base temperature K with incident energies meV and 2.63 meV, providing coverage of the entire first Brillouin zone at an elastic FWHM energy resolution of eV and 45 eV respectively. The resulting spectra were analyzed using the Horace [ewingsHORACESoftwareAnalysis2016] and Sunny [sunny_git] software packages.
@@ III.1 Magnetometry
| (K) | () | |
| -7.74 | 2.76 | |
| -7.98 | 2.80 | |
| -5.72 | 3.71 |
The bulk magnetic susceptibility along different crystallographic axes is shown in Figure 3(a). There are no signs of magnetic order down to 2 K and a simple fit to the inverse susceptibilities in the range K results in an average Weiss temperature K (see Table 2 for details), in agreement with previous reports [ashtarNewFamilyDisorderFree2020]. As the temperature is lowered, a pronounced easy axis anisotropy emerges along the -direction, while the planar response remains nearly axially symmetric over the entire temperature range. This should not be related directly to the single-ion anisotropy of Pr3+ though, as the bulk susceptibility measured in experiment represents an average over the six ions in the unit cell, whose easy axes are related by rotations in the hexagonal plane (see Figure 4(c)). In the low temperature limit, the susceptibility exhibits a small downturn, which could result either from interactions or from a saturation of the large Van-Vleck susceptibility expected for a singlet ground state system.
@@ III.2 CEF Spectrum
The inelastic neutron spectra measured on 11B-substituted powders of Pr3BWO9 are shown in Figure 4(a). Apart from the strong quasielastic scattering, the only prominent feature is an intense peak at meV. This mode nicely matches the energy scale seen in magnetometry. It softens slightly with increasing temperature due to thermal expansion and its integrated intensity closely follows the Debye-Waller factor, confirming it is of magnetic origin. Although additional features are observed at higher energies meV (not shown), they are all shown to be either spurious scattering in the monochromator/analyzer or phonons.
@@ III.3 Single-Ion Ground State
where are the Stevens operators [stevensMatrixElementsOperator1952, hutchingsPointChargeCalculationsEnergy1964] and is the free-ion Landé factor. However, due to the trivial magnetic point symmetry, there are 27 allowed crystal field parameters , making an unrestricted fit of experimental data to this Hamiltonian severely underconstrained. Instead, we rely on an effective point-charge model [scheiePyCrystalFieldSoftwareCalculation2021, dunEffectivePointchargeAnalysis2021, hutchingsPointChargeCalculationsEnergy1964]: An estimate of the CEF parameters is obtained directly from the experimentally determined crystal structure (see Sec. III.9) by calculating the electrostatic repulsion of neighboring ligands. We can diagonalize the resulting Hamiltonian to obtain the full spectrum and single-ion wavefunctions, allowing a direct comparison with experiment. By varying the effective charges of the three inequivalent O2- positions as well as one mean field exchange parameter in a simultaneous fit to INS, Magnetization and Susceptibility data, we obtain an accurate approximation of the single-ion Hamiltonian that reproduces all experimental results.
@@
The local frame of each Pr3+ ion is tilted by from the crystallographic -axis (see Figure 1(b)). In this reference frame, the quasi-doublet wavefunctions are well approximated by the (anti-)symmetric superpositions
The single-ion properties are axially symmetric, with perfect Ising anisotropy and due to the non-Kramers nature, while over 95% of the weights are concentrated on the maximal components. The weakened anisotropy observed in bulk arises naturally from averaging over the 6 equivalent Pr3+ ions in the unit cell. A sketch of the ground state anisotropy is depicted in Figure 4(c), illustrating the probability to observe a spin with any given orientation. Further details about the effective point charge model can be found in Appendix. A.